# A Bayesian approach to functional regression: theory and computation

### v2

@@ 3.1 Consistency in our RKHS model
The second conclusion is of certain relevance in itself, because the estimation of the number of components in mixture-like models is a hard problem in general (see Miller and Harrison, 2018, and references therein). Moreover, it is worth pointing out that the proof of Theorem 4 can be easily adjusted to guarantee consistency when the number of components is fixed beforehand (and thus the parameter space is finite-dimensional). Indeed, in this case Doob’s theorem applies directly under the sole condition that the times be distinct with prior probability one. The coefficients $\beta_{j}$ do not cause a problem for identifiability now, since the dimension of every parameter is the same.
@@
Assume model (3.3) in which the value of $p$ is fixed, and hence $\Theta_{p}$ is the (finite-dimensional) parameter space. Suppose that Condition 3-(i) holds as well. Then, if $\theta_{0}$ is the true value of the parameter, the posterior is consistent at $\theta_{0}$ with $\Pi$-probability one.
Note that by allowing $\beta_{j}$ to be zero in this situation we can sometimes circumvent the fact that the true value of the parameter might not have exactly $p$ components, as long as $p$ is larger than the true value $p(\theta_{0})$. Indeed, if $\theta_{0}\in\Theta$ with $p(\theta_{0})<p$ and $(X,Y)_{1:\infty}\sim P_{\theta_{0}}(X,Y)$ i.i.d., then we can find $\theta_{1}\in\Theta_{p}$, which is just $\theta_{0}$ completed with zeros, such that $P_{\theta_{0}}$(X, Y) = $P_{\theta_{1}}(X,Y)$ and the result holds almost surely.
A remark on Lebesgue consistency
All in all, Theorem 4 guarantees consistency for $\Pi$-almost every parameter in the support of the prior distribution. However, even though we can choose the prior so that $\operatorname{supp}(\Pi)=\Theta$, in principle there is no assurance that the $\Pi$-null set in which consistency may fail will not be a large set with respect to other (e.g. Lebesgue) measures. In fact, when the parameter space is infinite-dimensional there are examples of big inconsistency sets, even for reasonably chosen prior distributions (Diaconis and Freedman, 1986). Nonetheless, this problem can be alleviated when the parameter space is a countable union of disjoint finite-dimensional sets, as we can further refine our almost sure statement. First, note that there is a natural extension of the Lebesgue measure to our parameter space $\Theta$: just consider the genuine Lebesgue measure $\lambda_{p}$ on $\Theta_{p}$, and for all $B\subseteq\Theta$ measurable define $\lambda(B)=\sum_{p=1}^{\infty}\lambda_{p}(\Theta_{p}\cap B)$. Then, if we choose a prior distribution with respect to which this measure is absolutely continuous, the consistency set $\Theta_{*}$ in Theorem 4 will satisfy $\lambda(\Theta\setminus\Theta_{*})=0$. A similar approach is considered in Nobile (1994) and Miller (2023) to establish “Lebesgue”-almost sure consistency in finite mixture models with a prior on the number of components. In our case, the requirement of absolute continuity can be relaxed so that sets with nonzero Lebesgue measure have nonzero prior probability for some permutation of the component labels. We give more details on this result in Appendix B.
3.2 A sketch of the proof
We now present a proof of our consistency result. The general strategy will be to apply Doob’s theorem to a subset of the parameter space $\Theta$, where permutations are suitably taken into account and full identifiability holds, and then extend the conclusions to the whole parameter space. As we will see shortly, save for an eventual permutation, identifiability of the map $\theta\mapsto P_{\theta}(X,Y)$ is obtained in the RKHS case when the covariance function $K$ of the underlying stochastic process is non-degenerate.
Reduced parameter space. Consider the space $\mathbb{R}_{\delta}=(-\infty,-\delta]\cup[\delta,+\infty)$, with $\delta$ the fixed tolerance given in Condition 3-(ii), and define
where $[0,1]^{p}_{\text{ord}}=\{(t_{1},\dots,t_{p})\in[0,1]^{p}:t_{1}<\cdots<t_{p}\}$. Now consider $\tilde{\Theta}=\bigcup_{p\in\mathbb{N}}\tilde{\Theta}_{p}$, and note that the sets $\tilde{\Theta}_{1},\tilde{\Theta}_{2},\ldots$ are still disjoint. Then, again by Proposition A.1 in Miller (2023), we can conclude that $\tilde{\Theta}$ is a complete separable metric space under the metric $d$ in (3.2). We will henceforth say that a parameter in $\tilde{\Theta}$ is “ordered”.
Transformation into ordered form. For $\theta\in\Theta_{p}$, define $T(\theta)=\theta[\nu]$ if there is a $\nu\in S_{p}$ such that $\theta[\nu]\in\tilde{\Theta}_{p}$; otherwise set $T(\theta)=\theta$. Note that we can only transform $\theta$ to be in $\tilde{\Theta}_{p}$ if the times $t_{j}$ are all distinct and the coefficients $\beta_{j}$ satisfy $|\beta_{j}|\geq\delta$. But since the prior distribution assigns probability one to both events by Condition 3, we have $\Pi(T(\bm{\theta})\in\tilde{\Theta})=1$. It will be useful later to observe that for any set $B\subseteq\tilde{\Theta}_{p}$, if we denote $B[\nu]=\{\theta[\nu]:\theta\in B\}$, we have
$\bigcup_{\nu\in S_{p}}B[\nu]=T^{-1}(B).$ (3.4)
Collapsed model. Note that $P_{T(\theta)}(X,Y)=P_{\theta}(X,Y)$, and let $\tilde{Q}$ denote the distribution of $T(\bm{\theta})$ restricted to $\tilde{\Theta}$. Then the following model holds on the reduced space $\tilde{\Theta}$:
$T(\bm{\theta})\sim\tilde{Q},$ (3.5)
[…]
$\alpha_{0}+\sum_{j=1}^{p(\theta)}\beta_{j}X(t_{j})=\alpha_{0}^{\prime}+\sum_{j=1}^{p(\theta^{\prime})}\beta^{\prime}_{j}X(t^{\prime}_{j}),$
where $t_{i}\neq t_{j}$ and $t_{i}^{\prime}\neq t_{j}^{\prime}$ for $i\neq j$. Reordering the terms and combining those where the impact points coincide, this means that a linear combination of marginals of the process $X$ equals a constant. By taking variances, we can see that all the coefficients in this linear combination must vanish, since the covariance function of the process is strictly positive definite. But $\beta_{j}$ and $\beta_{j}^{\prime}$ cannot be 00 for any $j$ (by definition of $\tilde{\Theta}$), so it must be the case that $\theta^{\prime}=\theta[\nu]$ for some $\nu\in S_{p}$, where $p=p(\theta)=p(\theta^{\prime})$. However, $t_{1}<\cdots<t_{p}$ and $t^{\prime}_{1}<\cdots<t^{\prime}_{p}$, so $\nu$ must be the identity permutation, that is, $\theta^{\prime}=\theta$, contradicting the initial assumption.
Applying Doob’s theorem. Next we analyze the conclusions of Doob’s theorem applied to the collapsed model (3.5): there exists $\tilde{\Theta}_{*}\subseteq\tilde{\Theta}$ with $\Pi(T(\bm{\theta})\in\tilde{\Theta}_{*})=1$ such that, if $T(\theta_{0})\in\tilde{\Theta}_{*}$ and it holds that $(X,Y)_{1:\infty}\sim P_{T(\theta_{0})}(X,Y)$ i.i.d., then for any neighborhood $B\subseteq\tilde{\Theta}$ of $T(\theta_{0})$ we have
[…]
Now define $\Theta_{*}$ to be the set of all parameters in $\Theta$ that can be obtained by permuting a parameter in $\tilde{\Theta}_{*}$, i.e., $\Theta_{*}=\bigcup_{p=1}^{\infty}\bigcup_{\nu\in S_{p}}(\tilde{\Theta}_{*}\cap\tilde{\Theta}_{p})[\nu]$. Then, by (3.4) we get
$\Pi(\bm{\theta}\in\Theta_{*})=\Pi\left(T(\bm{\theta})\in\bigcup_{p=1}^{\infty}(\tilde{\Theta}_{*}\cap\tilde{\Theta}_{p})\right)=\Pi(T(\bm{\theta})\in\tilde{\Theta}_{*})=1.$
$B=\left\{\theta\in\tilde{\Theta}:d(\theta,T(\theta_{0}))<\varepsilon\right\}.$ (3.7)
Observe that, since $\varepsilon<1$, we have $B\subseteq\tilde{\Theta}_{p_{0}}$ by definition of $d$. Moreover, $\bigcup_{\nu\in S_{0}}B[\nu]\subseteq\tilde{B}(\theta_{0},\varepsilon)$. Then, again by (3.4), we can write
$\Pi(\bm{\theta}\in\tilde{B}(\theta_{0},\varepsilon)\mid(X,Y)_{1:n})$ $\geq\Pi(\bm{\theta}\in\bigcup_{\nu\in S_{0}}B[\nu]\mid(X,Y)_{1:n})$ (3.8)
$=\Pi(T(\bm{\theta})\in B\mid(X,Y)_{1:n}).$
[…]
$\Pi(\mathcal{P}=p_{0}\mid(X,Y)_{1:n})$ $=\Pi(\bm{\theta}\in\Theta_{p_{0}}\mid(X,Y)_{1:n})$
$\geq\Pi(\bm{\theta}\in\tilde{B}(\theta_{0},\varepsilon)\mid(X,Y)_{1:n})$
[…]
As a final comment, it is worth reiterating that in this theoretical aspect of our work we have closely followed the techniques recently developed in Miller (2023), where the author provides a simplification of the work by Nobile (1994) in studying posterior consistency in finite-dimensional mixture models with a prior on the number of components. While the methods are quite similar, we have succeeded in extending this theory to a fundamentally different situation, namely functional regression, which thanks to the RKHS formulation shares the key properties that allow for a treatment analogous to the finite-dimensional mixture case.
@@ 4 Experimental results
In this section we present the results of the experiments carried out to test the performance of our Bayesian methods in different scenarios, together with an overview of their computational implementation. Further details such as simulation parameters or algorithmic decisions, as well as additional experiments, figures and tables are provided in Appendices AC and D, while the code itself is publicly available on GitHub at https://github.com/antcc/rk-bfr-jump.
For simulated data we consider $n=200$ training samples and $n^{\prime}=100$ testing samples on an equispaced grid of $100$ points on $[0,1]$, and for the real data sets we perform a 66%/33% train/test split. We fit our models on the training data, using RJMCMC to approximate the posterior, and then use the test set for out-of-sample predictions. We independently repeat the whole process 10 times (each with a different train/test configuration) to account for the stochasticity in the sampling step, and average the results across these executions. For the purposes of prediction we consider the point statistics trimmed mean (10%), median and mode to aggregate together predictions and summarize parameters (see Section 2.3). Moreover, the values of $\{t_{j}\}$ that fall outside the specified grid are truncated to the nearest neighbor in the grid, with the additional restriction that time instants in different components of our models cannot be equal.
The Python library used to perform the RJMCMC approximation is Eryn (Karnesis et al., 2023). It is a toolbox for Bayesian inference that allows trans-dimensional posterior approximation, running multiple chains in an ensemble configuration with different starting points, and incorporating a parallel tempering mechanism (Hukushima and Nemoto, 1996) to increase both convergence speed and the acceptance rate. Because of execution time constraints, the hyperparameters of the Eryn sampler are selected manually based on an initial set of experiments, as well as recommendations from the authors. Moreover, some amount of post-processing is needed to mitigate the well-known label switching phenomenon that occurs in MCMC approximations of mixture-like models (Stephens, 2000); see Appendix A.2.
@@
We employ two different types of simulated data sets, all with a common value of $\alpha_{0}=5$ and $\sigma^{2}=0.5$:
@@
Again we consider two different types of simulated data sets, with a common value of $\alpha_{0}=-0.5$:
@@
Additionally, we use three real data sets well known in the literature. The first one is a subset of the Medflies data set (Carey et al., 1998), consisting on samples of the number of eggs laid daily by 534 flies over 30 days, to predict whether their longevity is high or low. The second one is the Berkeley Growth Study data set (Tuddenham and Snyder, 1954), which records the height of 54 girls and 39 boys over 31 different points in their lives. Finally, we selected a subset of the Phoneme data set (Hastie et al., 1995), based on 200 digitized speech frames over 128 equispaced points to predict the phonemes “aa” and “ao”.
@@
We have adopted a visual approach to presenting the experimentation results, using colored graphs to help visualize them. In each case, the mean and standard deviation of the score obtained across the 10 random runs is shown, depicting our methods in blue and the comparison algorithms in orange. We also show the global mean of all the comparison algorithms with a dashed vertical line, excluding extreme negative results to avoid distortion. Moreover, we separate one-stage and two-stage methods, the latter being the ones that perform variable selection or dimension reduction prior to a multiple linear/logistic regression method (represented with “+r”/“+log” in the figures). We name our methods according to the acronyms described in Section 2.3.

### v3

@@ 3.1 Consistency via Doob’s theorem
The hypothesis of positive definiteness of $K$ is needed to ensure identifiability. On the other hand, the second conclusion is of certain relevance in itself, because the estimation of the number of components in mixture-like models is a hard problem in general (see Miller and Harrison, 2018, and references therein). Moreover, it is worth pointing out that the proof of Theorem 5 can be easily adjusted to guarantee consistency when the number of components is fixed beforehand and the parameter space is finite-dimensional.
A remark on Lebesgue consistency
All in all, Theorem 5 guarantees consistency for $\Pi$-almost every parameter in the support of the prior distribution. However, even though we can choose the prior so that $\operatorname{supp}(\Pi)=\Theta$, in principle there is no assurance that the $\Pi$-null set in which consistency may fail will not be a large set with respect to other measures. In fact, when the parameter space is infinite-dimensional there are examples of big inconsistency sets, even for reasonably chosen prior distributions (Diaconis and Freedman, 1986). Nonetheless, this problem can be alleviated when the parameter space is a countable union of disjoint finite-dimensional sets. First, note that there is a natural extension of the Lebesgue measure to our parameter space $\Theta$: just consider the genuine Lebesgue measure $\lambda_{p}$ on $\Theta_{p}$, and for all $B\subseteq\Theta$ measurable define $\lambda_{\infty}(B)=\sum_{p=1}^{\infty}\lambda_{p}(\Theta_{p}\cap B)$. Then, if we choose a prior distribution with respect to which this measure is absolutely continuous, the inconsistency set in Theorem 5 will satisfy $\lambda_{\infty}(\Theta\setminus\Theta_{*})=0$ and thus be “small” with respect to a Lebesgue-type measure. In our case, the requirement of absolute continuity can be relaxed so that sets with nonzero Lebesgue measure have nonzero prior probability for some permutation of the component labels.
Proposition 6.
Suppose that Condition 4 holds. Furthermore, assume that for all $p\in\mathbb{N}$ we have
  1. (i)
    $\Pi(\mathcal{P}=p)>0$.
  2. (ii)
    $\sum_{\nu\in S_{p}}\Pi(\theta[\nu]\in B|\mathcal{P}=p)=0$ implies $\lambda_{p}(B)=0$, for all $B\subseteq\Theta_{p}$ measurable.
Then the conclusion of Theorem 5 remains valid with $\lambda_{\infty}(\Theta\setminus\Theta_{*})=0$.
The first condition is a somewhat technical requirement. The second condition is met, for example, if $\theta|p$ has a density with respect to the Lebesgue measure that is invariant to permutations of the component labels and positive on all of $\Theta_{p}$. A similar approach is considered in Nobile (1994) and Miller (2023) to establish “Lebesgue”-almost sure consistency in finite mixture models with a prior on the number of components.
3.2 Consistency and contraction rates via Schwartz’s theorem
The Doob-type results of the previous section, although interesting, are still assertions on consistency from the point of view of the prior distribution. We now seek additional results that offer more definitive statements and, more importantly, establish posterior contraction rates. In a nonparametric setting where the object of interest is a probability density, there is a stronger consistency result by Schwartz (1965) which omits the $\Pi$-almost sure qualification under some more restrictive conditions, though it requires a dominated model and a density to estimate. To this end, we consider the conditional model $Y|X=x,\theta\sim f_{\theta}(y|x)$, where $\theta\in\Theta$ is a parameter vector and $f_{\theta}(y|x)$ is the density of $Y$ given $X=x$ with respect to the Lebesgue measure $\lambda$ on $\mathcal{Y}$ (which under our assumptions is the normal density given by (2.2)). We could work with this model and analyze the fixed design case with non-i.i.d. data (see Choi and Ramamoorthi, 2008), but we choose to focus on the arguably more significant random design case.
In general, in infinite-dimensional models there is no reference measure with respect to which to define a density. However, in our specific functional regression setting we can find such a density through the following observation: it follows from the disintegration theorem and a straightforward application of Dynkin’s $\pi$-$\lambda$ theorem that the function $(x,y)\mapsto f_{\theta}(y|x)$ is a joint density of $(X,Y)$ with respect to the product measure $\rho=Q_{X}\times\lambda$, where $Q_{X}$ is the law of the process $X$. In this way, denoting $f_{\theta}:=f_{\theta}(y|x)$, we can express our full model (with respect to the product measure $\rho$) as
$(X,Y)_{1:n}\mid f_{\theta}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}f_{\theta}\quad\text{and}\quad f_{\theta}\sim\Pi_{\mathcal{F}},$
where $\Pi_{\mathcal{F}}$ is a prior distribution on the parameter class $\mathcal{F}=\{f_{\theta}:\theta\in\Theta\}$, which is the set of all densities following our model relative to the dominating measure $\rho$ on $\mathcal{X}\times\mathcal{Y}$. In practice, we work with the prior distribution $\Pi$ on $\Theta$ specified in (3.2), and interpret $\Pi_{\mathcal{F}}$ as the pushforward measure of $\Pi$ by the measurable mapping $\Phi:\Theta\to\mathcal{F}$ given by $\Phi(\theta)=f_{\theta}$. As always, we assume the existence of a “true” density $f_{0}:=f_{\theta_{0}}(y|x)\in\mathcal{F}$ that generates the observations, and denote $\Theta_{0}\equiv\Theta_{p(\theta_{0})}$.
Now we introduce a few concepts that lie at the core of most extensions of Schwartz’s consistency theorem. Recall that the Kullback-Leibler divergence from $f_{0}$ to $f_{\theta}$ is defined as $D_{\mathrm{KL}}(f_{0}\,\|\,f_{\theta})=\int f_{0}\log(f_{0}/f_{\theta})\,d\rho$. We say that $f_{0}$ belongs to the Kullback-Leibler support of $\Pi_{\mathcal{F}}$, and write it as $f_{0}\in\operatorname{KL}(\Pi_{\mathcal{F}})$, if $\Pi_{\mathcal{F}}\left(f_{\theta}:D_{\mathrm{KL}}(f_{0}\,\|\,f_{\theta})<\varepsilon\right)>0$ for all $\varepsilon>0$. This can be interpreted as saying that the prior places sufficient mass “near” $f_{0}$. On the other hand, let $N(\varepsilon,\mathcal{F},d)$ denote the $\varepsilon$-covering number of the set $\mathcal{F}$ with respect to the distance $d$, i.e., the minimal number of $d$-balls of radius $\varepsilon$ needed to cover $\mathcal{F}$. Finally, define the Hellinger distance $d_{H}$ between two densities as $d^{2}_{H}(f_{\theta},f_{\theta^{\prime}})=1-\int\sqrt{f_{\theta}f_{\theta^{\prime}}}\,d\rho$. The following result will give us consistency in the sense of Schwartz, by means of a sieve that controls the complexity of the model as the sample size $n$ increases, measured in terms of the metric entropy (the logarithm of the $\varepsilon$-covering number) of the model.
Theorem 7 (Theorem 6.23 in Ghosal and van der Vaart (2017)).
Suppose that for every $\varepsilon>0$ there exist measurable sets $\mathcal{F}_{n}\subseteq\mathcal{F}$ and a constant $C>0$ such that, for sufficiently large $n$,
  1. (i)
    $\log N(\varepsilon,\mathcal{F}_{n},d_{H})\leq n\varepsilon^{2}$.
  2. (ii)
    $\Pi_{\mathcal{F}}(\mathcal{F}\setminus\mathcal{F}_{n})\leq\exp\{-Cn\}$.
Then the posterior distribution is consistent relative to $d_{H}$ at every $f_{0}\in\operatorname{KL}(\Pi_{\mathcal{F}})$.
Looking at the form of the parameter space $\Theta$ in (3.1), a natural choice for the sieve is $\Theta_{n}=\bigcup_{p=1}^{p_{n}}\Theta_{p}$, for $p_{n}=O(h(n))$ and $h(n)\to\infty$ a function to be determined later, with the corresponding sieve of densities defined as $\mathcal{F}_{n}=\{f_{\theta}\in\mathcal{F}:\theta\in\Theta_{n}\}$. Moreover, to keep the conditions as simple as possible, it is more convenient to work with compact parameter spaces $\Theta_{p}$ and adapt the prior distributions in (3.2) accordingly; we assume this is so for the remainder of the section. Plus, we need to impose some smoothness restrictions on the underlying process $X$.
Condition 8 (Functional constraints).
Suppose that each trajectory $x$ of $X$ is $Q_{X}$-almost surely Lipschitz with Lipschitz constant $L(x)>0$, and assume that $\mathbb{E}_{Q_{X}}[L^{2}]<\infty$.
See Appendix B for a brief discussion and examples of processes that satisfy this condition. We can now state a new Hellinger consistency result for our model, which transforms the requirement that $f_{0}\in\operatorname{KL}(\Pi_{\mathcal{F}})$ into a condition on the prior $\Pi$ on $\Theta_{0}$ and places an additional mild constraint on the tails of the marginal prior distribution of $p$. These conditions guarantee that we can choose a suitable growth order for $p_{n}$ in the sieve to apply Theorem 7.
Theorem 9.
Assume Condition 8 holds, and suppose that $\Pi(p)=O(e^{-\delta p(\log p)^{k}})$ as $p\to\infty$, with $\delta>0$ and $k>1$. If $\Pi$ assigns positive mass to every open neighborhood of $\theta_{0}$ in $\Theta_{0}$, then the posterior distribution in the model $\theta\sim\Pi$ and $(X,Y)_{1:n}|\theta\sim f_{\theta}$ i.i.d. is consistent relative to $d_{H}$ at $f_{0}$, i.e., for every $\varepsilon>0$ it holds that
$\lim_{n\to\infty}\Pi_{n}(\theta:d_{H}(f_{0},f_{\theta})<\varepsilon\mid(X,Y)_{1:n})=1\quad P_{0}^{\infty}(X,Y)-\text{a.s.}$
An immediate consequence is that if $\Pi(p)$ is supported on a finite set (as we do in our experiments), then the prior tail condition is automatically verified. Moreover, since it only needs to hold for large values of $p$, we can always consider a hybrid prior distribution that changes at some large cutoff point to have rapidly decaying tails.
Lastly, to find posterior contraction rates for our model we consider an extension of Theorem 7 that employs the so-called second Kullback-Leibler variation, defined by $V_{2}(f_{0},f_{\theta})=\mathbb{E}_{f_{0}}[(\log(f_{0}/f_{\theta})-D_{\mathrm{KL}}(f_{0}\,\|\,f_{\theta}))^{2}]$, to quantify the Kullback-Leibler property of $f_{0}$ in terms of convergence speed. In particular, for $\varepsilon>0$ we consider the neighborhood
$B_{2}(f_{0},\varepsilon)=\{f\in\mathcal{F}:D_{\mathrm{KL}}(f_{0}\,\|\,f_{\theta})<\varepsilon^{2},\ V_{2}(f_{0},f_{\theta})<\varepsilon^{2}\}.$
Theorem 10 (Theorem 8.9 in Ghosal and van der Vaart (2017)).
Suppose that there exist measurable sets $\mathcal{F}_{n}\subseteq\mathcal{F}$ and a constant $C>0$ such that, for a sequence $\varepsilon_{n}\to 0$ with $n\varepsilon_{n}^{2}\to\infty$, the following hold for sufficiently large $n$:
  1. (i)
    $\log N(\varepsilon_{n}/2,\mathcal{F}_{n},d_{H})\leq n\varepsilon_{n}^{2}$.
  2. (ii)
    $\Pi_{\mathcal{F}}(\mathcal{F}\setminus\mathcal{F}_{n})\leq\exp\{-(C+4)n\varepsilon_{n}^{2}\}$.
  3. (iii)
    $\Pi_{\mathcal{F}}(B_{2}(f_{0},\varepsilon_{n}))\geq\exp\{-Cn\varepsilon_{n}^{2}\}$.
Then the posterior rate of contraction at $f_{0}$ with respect to $d_{H}$ is $\varepsilon_{n}$.
Note that the condition $n\varepsilon_{n}^{2}\to\infty$ excludes the parametric rate $n^{-1/2}$, since in nonparametric situations the rates are usually slower. Condition (i) bounds the complexity of the model, condition (ii) merely expresses that subsets of negligible prior mass do not play a role in the rate of contraction, and condition (iii) ensures that sufficient prior mass is put “near” the true density $f_{0}$. Following a similar proof strategy as in Theorem 9, we can find a convenient sieve to apply Theorem 10 and prove the following result in our model.
Theorem 11.
Assume Condition 8 holds. Let $0<\gamma<1/2$ and suppose that the prior distribution satisfies $\Pi(p)=O(e^{-\delta p\log p})$ as $p\to\infty$, with $\delta>16(3-4\gamma)/(1-2\gamma)$. If $\Pi$ has a density on $\Theta_{0}$ with respect to Lebesgue measure that is bounded away from zero in a neighborhood of $\theta_{0}$, then the posterior distribution in the model $\theta\sim\Pi$ and $(X,Y)_{1:n}|\theta\sim f_{\theta}$ i.i.d. contracts at a rate $\varepsilon_{n}=n^{-\gamma}$ relative to $d_{H}$ at $f_{0}$, i.e., for every $M_{n}\to\infty$ it holds that
$\Pi_{n}(\theta:d_{H}(f_{0},f_{\theta})\geq M_{n}\varepsilon_{n}\mid(X,Y)_{1:n})\to 0\quad\text{in }P_{0}^{(n)}(X,Y)\text{-probability}.$
The requirements of this result are akin to those of Theorem 9, but the conclusion is considerably stronger. To verify the condition on the density of $\Pi$, provided that it exists, it suffices to show that it is continuous and positive at $\theta_{0}$ (which our proposed prior distributions satisfy). The contraction rate $\varepsilon_{n}=n^{-\gamma}$ with $0<\gamma<1/2$ is optimal in the sense that it is as close as desired to the parametric rate of $n^{-1/2}$, and it is minimax optimal for estimators of $f_{\theta}$ given the model $\mathcal{F}_{n}$ (see Ghosal and van der Vaart, 2017, p. 198 and references therein).
To conclude, it is worth pointing out that a contraction rate relative to $d_{H}$ is also a contraction rate relative to any distance bounded above by a multiple of the Hellinger distance, so we immediately have consistency and a contraction rate in more descriptive distances. Let $h]_{M}$ be the real-valued function $h$ truncated to the interval [M,M][-M,M].
@@ 3.2 Consistency and contraction rates via Schwartz’s theorem
Under the same conditions as in Theorem 11, $\varepsilon_{n}=n^{-\gamma}$ is a posterior contraction rate relative to the total variation distance $d_{1}(f_{0},f_{\theta})=\int|f_{0}-f_{\theta}|\,d\rho$ and the mean-variance discrepancy metric $d_{2,Q_{X}}(\theta_{0},\theta)=\|[\mu_{\theta_{0}}]_{M}-[\mu_{\theta}]_{M}\|_{2,Q_{X}}+|\sigma_{0}^{2}-\sigma^{2}|$, for any $M>0$, where $\|\cdot\|_{2,Q_{X}}$ is the $L^{2}(Q_{X})$-norm and $\mu_{\theta}(x)=\alpha+\sum_{j=1}^{p(\theta)}\beta_{j}x(t_{j})$.
@@ 4 Experimental results
In this section we present the results of the experiments carried out to test the performance of our Bayesian methods in different scenarios, together with an overview of their computational implementation. Further details such as simulation parameters, execution times or algorithmic decisions, as well as additional experiments, figures and tables are provided in Appendices A, C and D, while the code itself is publicly available on GitHub at https://github.com/antcc/rk-bfr-jump.
For simulated data we consider $n=200$ training samples and $n^{\prime}=100$ testing samples, with functional regressors measured on an equispaced grid of $100$ points on $[0,1]$, and for the real data sets we perform a 66%/33% train/test split. We fit our models on the training data, using RJMCMC to sample from the approximate posterior, and then compute out-of-sample predictions on the testing set. We independently repeat the whole process 10 times (each with a different train/test configuration) to account for the stochasticity in the sampling step, and average the results across these executions. For the purposes of prediction we consider the point statistics trimmed mean (10%), median and mode to aggregate together predictions and summarize parameters (see Section 2.3). Moreover, the values of $\{t_{j}\}$ that fall outside the specified grid are truncated to the nearest neighbor in the grid, with the additional restriction that time instants in different components of our models cannot be equal.
The Python library used to perform the RJMCMC approximation is Eryn (Karnesis et al., 2023). It is a toolbox for Bayesian inference that allows trans-dimensional posterior approximation, running multiple chains in an ensemble configuration with different starting points, and incorporating a parallel tempering mechanism (Hukushima and Nemoto, 1996) to increase both convergence speed and acceptance rate. Because of execution time constraints, the hyperparameters of the Eryn sampler are selected manually based on preliminary experiments and the authors’ recommendations. This reduction in computational cost renders the sampling step practically viable; information on the execution times is provided in Appendix C.4. Moreover, some amount of post-processing is needed to mitigate the well-known label switching phenomenon that occurs in MCMC approximations of mixture-like models (Stephens, 2000); see Appendix A.2.
@@
We employ two different types of simulated data sets, all with a common value of $\alpha=5$ for the intercept and $\sigma^{2}=0.5$ for the error variance:
@@
Again we consider two different types of simulated data sets, with a common value of $\alpha=-0.5$ for the intercept:
@@
Additionally, we use three real data sets well known in the literature. The first one is a subset of the Medflies data set (Carey et al., 1998), consisting on samples of the number of eggs laid daily by 534 flies over 30 days, to predict whether their longevity is high or low. The second one is the Berkeley Growth Study data set (Tuddenham and Snyder, 1954), which records the height of 54 girls and 39 boys over 31 different points in their lives. Finally, we selected a subset of the Phoneme data set (Hastie et al., 1995) based on 200 digitized speech frames over 128 equispaced points to predict the phonemes “aa” and “ao”.
@@
We have adopted a visual approach to presenting the experimentation results, using colored graphs to help visualize them. In each case, the mean and standard deviation of the score obtained across the 10 independent runs is shown, depicting our methods in blue and the comparison algorithms in orange. We also show the global mean of all the comparison algorithms with a dashed vertical line, excluding extreme negative results to avoid distortion. Moreover, we separate one-stage and two-stage methods, the latter being the ones that perform variable selection or dimension reduction prior to a multiple linear/logistic regression method (represented with “+r”/“+log” in the figures). We name our methods according to the acronyms described in Section 2.3.